13 Find all solutions to 6 tan” x 20 tan x + 6 = 0 14 Calculate and find the modulus of (3 – V2i)?
Answer #1

e2= 1) 1e=VIT r cos o 3 6tean²-20 tans +6=0 nove, let z = 3-v2i put tean x= a e cos o=3, rsino = -2 6a² – 20 a +6=0 (3a?- loa I hope you have got your answer , if you need any further clarification in any step, please let me know through comment.